OpenStudy (anonymous):

10ml of gaseous hydrocarbon were exploded with 70 ml(excess) of oxygen. after explosion the residual gases had a volume of 50ml which was reduced to20ml on exposure to alkali identify the hydrocarbon ( all volume measurements were made at spt)

OpenStudy (inkyvoyd):

Oh man.

OpenStudy (anonymous):

what i will give the options c2h4 c4h8 c3h8 c4h10

OpenStudy (anonymous):

the answer is c3h8 plz try to do it

OpenStudy (inkyvoyd):

Well, maybe we try to calculate the the molar quantities of these gasses.

OpenStudy (inkyvoyd):

c2h4:28 c4h8:56 c3h8:44 c4h10:58

OpenStudy (anonymous):

ok then......

OpenStudy (inkyvoyd):

No, I don't know. Uh, if you can wait mayb 10 or 20 mins, my mom is coming to pick me up, and she should know (she was a chem engineer)

OpenStudy (inkyvoyd):

The key here is probably to look at the hydrocarbon itself I bet.

OpenStudy (anonymous):

ok i will wait

OpenStudy (inkyvoyd):

Different hydrocarbons most likely have different burning reactions

OpenStudy (anonymous):

any hydrocarbons when burned will give out c02 h20 and heat

OpenStudy (inkyvoyd):

Why does volume reduce on exposure to alkali?

OpenStudy (anonymous):

i am also confused

OpenStudy (anonymous):

plzz dont forget to ask ur mother>>

OpenStudy (inkyvoyd):

She's coming right now :)

OpenStudy (inkyvoyd):

Hopefully she didn't forget.

OpenStudy (anonymous):

ya u r right

OpenStudy (inkyvoyd):

uh, 50ml is CO2

OpenStudy (anonymous):

yes

OpenStudy (inkyvoyd):

plus O2 actually. My mom's not helping -.- . Here, let me go home, and hten bother her more lol.

OpenStudy (inkyvoyd):

cya in 5, maybe 10 mins.

OpenStudy (anonymous):

ok

OpenStudy (inkyvoyd):

Let me make a note before I leave, so I don't forget myself CnHm+(m/4)O2=nCO2+XO2

OpenStudy (inkyvoyd):

I have it @shameer1 !

OpenStudy (anonymous):

what is n , m and x here plzzz explain

OpenStudy (inkyvoyd):

the residual gases had a volume of 50ml which was reduced to20ml on exposure to alkali basically, there residual gas currently has oxygen and carbon dioxide; when exposed to alkali, the carbon dioxide reacts with the alkali to make salt, and what remains is oxygen. In any case what is left is 20 ml of oxygen.

OpenStudy (inkyvoyd):

in other words, we can determine that 10 ml of the gaseous hydrocarbon reacted with 70-20ml of oyxgen, or 50 ml of oxygen.

OpenStudy (inkyvoyd):

CnHm+(n+m/4)O2=nCO2+(m/2)H2O ->n and m are unknown numbers. CnHm+(n+m/4)O2=nCO2+(m/2)H2O we know that we have 10 ml of hydrocarbon, and 50 ml of oxygen (this is how much reacted), and we got 30 ml of CO2

OpenStudy (anonymous):

what is the final answer u get

OpenStudy (inkyvoyd):

not done :S

OpenStudy (inkyvoyd):

Find the vapor density of the selections first c2h4:28 c4h8:56 c3h8:44 c4h10:58

OpenStudy (inkyvoyd):

@shameer1 ?

OpenStudy (anonymous):

then what to do

OpenStudy (inkyvoyd):

Well, we need to find the ratio of c to h

OpenStudy (inkyvoyd):

@FoolForMath fool fool, helpy helpy

OpenStudy (anonymous):

plz can u do it and show me

OpenStudy (anonymous):

did ur mother help u

OpenStudy (inkyvoyd):

No, she short of ignored me, lol. I'm going to tag some more people, because I just realized that I might have to go soon, and I'm not sure if I'm even doing the problem correctly. @Diyadiya @nbouscal @KingGeorge

OpenStudy (anonymous):

guys do u heve idea to solve this

OpenStudy (anonymous):

no sorry shameer i tried but dunno :(

OpenStudy (anonymous):

oh...no..no

OpenStudy (anonymous):

@JFraser @shankvee

OpenStudy (inkyvoyd):

I'm going to repeat the problem, because it is way too far up. I will also add what I know. 10ml of gaseous hydrocarbon were exploded with 70 ml(excess) of oxygen. after explosion the residual gases had a volume of 50ml which was reduced to20ml on exposure to alkali identify the hydrocarbon ( all volume measurements were made at spt) We know that: 20 ml is oxygen, and 30 ml is CO2 (residual gas)

OpenStudy (inkyvoyd):

CH being the unknown hydrocarbon.

OpenStudy (inkyvoyd):

CH (10 ml)+O2 (50 ml)->[CO2](30ml)+H2O

OpenStudy (inkyvoyd):

what does the molecular structure of a gas have to do with its density @shameer1

OpenStudy (anonymous):

molecular mass = 2* VD

OpenStudy (inkyvoyd):

any idea what we can do with it and the equation?

OpenStudy (inkyvoyd):

VD of oxygen would be 16. VD of CO2?

OpenStudy (vincent-lyon.fr):

The answer is C3H8. Ok! Balance the equation and do the problem the 'forward-way' (knowing 10mL of C3H8 and finding the 30mL CO2 and 20mL excess O2). You will see the mechanism of it, and then, you can solve your problem, which is worded the 'reversed-way'.

OpenStudy (anonymous):

we have to find the formula of hydrocarbon

OpenStudy (mos1635):

@Vincent-Lyon.Fr why we don have to indlude vapours of H2O to gases??

OpenStudy (vincent-lyon.fr):

In standard conditions, water is liquid, so its volume is negligible.

OpenStudy (mos1635):

you are so right AGAIN!!!!!

OpenStudy (mos1635):

all together CnHm + (4n+m)/4 O2 --> nCO2 + (m/2)H2O initial 10 70 0 0 reacts/produce -10 -10*(4n+m)/4 +10n +10*(m/2) final 0 70-10(4n+m)/4 10n 10*(m/2) by given data you have: Vco2=30 10n =30 n=3 and Vo2=20 (H20 is liquid in STP) 70-10(4n+m)/4 =20 (n=3) solve to m .....m=8 so C3H8

OpenStudy (vincent-lyon.fr):

@shameer1: have you tried doing what I suggested?

OpenStudy (vincent-lyon.fr):

Try to solve that problem: "10 mL of C3H8 are mixed with 70mL of O2. What is the residual gas volume after explosion? An alkali sucks in all the CO2 formed. What is the volume of gas left?" Then you will see how it all works and you will be able to solve the initial problem, because you will understand the inner-mechanism behind it.

OpenStudy (anonymous):

honestly i dint read all replies but did you manage to solve the problem?

OpenStudy (anonymous):

can some one do this question in other way i cant understand those which are witten plzzz

OpenStudy (paxpolaris):

for constant temperature & pressure: number of moles is proportional to volume. @shameer1 can't show any different way to do this follow mos1635 or ask what you don't understand

OpenStudy (anonymous):

Vco2=30 how the volume became 30

OpenStudy (paxpolaris):

CO2 is produced in reaction ... but then it reacts with alkali to produce salt (since CO2 is acidic). So, Vol. CO2 = 50 - 20 = 30ml

OpenStudy (anonymous):

what is this 70-10(4n+m)/4

OpenStudy (anonymous):

plz explain this paxx

OpenStudy (paxpolaris):

70 is initial volume of O2 minus the vol.of O2 used \(10 \times {\left( 4m+n \right)}/ 4\) is equal to the final volume of O2 i.e. 20

OpenStudy (paxpolaris):

\[\Large C_3H_8+5O_2 \rightarrow 3CO_2+4H2O\]

OpenStudy (paxpolaris):

70 is initial volume of O2 minus the vol.of O2 used 10×(4m+n)/4 is equal to the excess volume of O2 i.e. 20